If a particle of charge 10^-12C moving with velocity of 10^5 m/s experience a force of 10^-10 N in y direction due to magnetic field,then the minimum value of magnetic field is?

Grade:12

1 Answers

Arun
6 years ago
Dear Rajit
By Lorentz Force
F = q * v* B sin@
Now
B = F/q*v* sin@
For minimum magnetic field sin@ is maximum = 1
Hence
B = 10-10/105 *10-12 
B = 10-3 T in Z direction
 
Regards
Arun (askIITians forum expert)